# SMReferences for Appendix

### v3

@@ 4.2 Extension to approximately sparse GLMs
In previous sections, we have assumed that at least one of the nuisance parameters, $\pi^{\ast}$ or $r^{\ast}$, exactly follows a sparse GLM, with sparsity below $\sqrt{n}/\log p$. Belloni et al. (2014); Chernozhukov et al. (2018); Bradic et al. (2019a) and Smucler et al. (2019) considered the so-called approximately sparse GLMs to model the nuisance parameters. Similar to Condition 3, the condition below imposes that at least one of the two nuisance parameters, $r^{\ast}$ or $\pi^{\ast}$, follows the approximately sparse GLM.
@@
There exist two (nonlinear) monotonically increasing, twice-differentiable link functions $\phi,\psi$ with uniformly bounded first and second derivatives, such that the following hold: for any positive integer $k$
  1. (i)
    The OR model can be approximated by a GLM with (nonlinear) link $\psi$ using $s$ covariates with approximation error decaying with rate $s^{-\xi_{r}}$ for some $\xi_{r}>0$: there exists $\beta_{s}^{\ast}$ such that $\|\beta_{s}^{\ast}\|_{0}\leq s$ and $\mathbb{E}[(r^{\ast}(X)-\psi(\beta_{s}^{\ast\top}X))^{2}]\ll s^{-2\xi_{r}}$;
  2. (ii)
    The PS model can be approximated by a GLM with (nonlinear) link $\psi$ using $s$ covariates with approximation error decaying with rate $s^{-\xi_{r}}$ for some $\xi_{\pi}>0$: there exists $\gamma_{s}^{\ast}$ such that $\|\gamma_{s}^{\ast}\|_{0}\leq s$ and $\mathbb{E}[(\pi^{\ast}(X)-\phi(\gamma_{s}^{\ast\top}X))^{2}]\ll s^{-2\xi_{\pi}}$.
We silence the dependence on $s$ by denoting $\beta^{\ast}$ and $\gamma^{\ast}$ when $s=p$; in particular, without loss of generality, we take $\beta^{\ast}$ and $\gamma^{\ast}$ to minimize $\mathbb{E}[(r^{\ast}(X)-\psi(\beta^{\top}X))^{2}]$ and $\mathbb{E}[(\pi^{\ast}(X)-\phi(\gamma^{\top}X))^{2}]$. Bradic et al. (2019a) proved the following necessary condition for the existence of $\sqrt{n}$-consistent estimator of the ATE $\tau$ when $p\gtrsim n$:
$\xi_{r}\vee\xi_{\pi}>1/2\text{ by choosing $s\asymp\sqrt{n/\log p}$}.$ (17)
For simplicity, we take $p\asymp n^{2}$. Although it is possible to relax the lower bound requirement on $p$ by imposing conditions similar to Assumption 9 of Bradic et al. (2019a), we decide not to further pursue this in this paper. As a corollary of the results in Section 3, the DCal estimator achieves $\sqrt{n}$-consistency under (17). To state this result, we need to modify the regularity conditions on the OR and PS coefficient estimates computed from the training dataset $\mathcal{D}_{\mathrm{tr}}$ as follows. First, we need to modify Condition 7 as
@@
Condition 11(i) holds true and there exists a constant $m_{\psi}$ such that for any $t$, $\psi^{\prime}(t),\psi^{\prime\prime}(t)\leq m_{\psi}$. Moreover, when $\xi_{r}>1/2$ and $\xi_{\pi}>0$, with probability converging to $1$, there exists $\hat{\beta}$ computed from $\mathcal{D}_{\mathrm{tr}}$ such that
We then need to modify Condition 8 as
Condition 13.
Condition 11(ii) holds true and there exists a constant $m_{\phi}$ such that for any $t$, $\phi^{\prime}(t),\phi^{\prime\prime}(t)\leq m_{\phi}$. Moreover, when $\xi_{\pi}>1/2$ and $\xi_{r}>0$, with probability converging to $1$, there exists $\hat{\gamma}$ computed from $\mathcal{D}_{\mathrm{tr}}$ such that
@@
The existence of $\hat{\beta}$ and $\hat{\gamma}$ satisfying Condition 12 and Condition 13 can follow similarly from the proof in Appendix B of Bradic et al. (2019a) or from Smucler et al. (2019). In fact, Condition 12 and Condition 13 closely resemble Condition 7 and Condition 8, respectively, which are standard results of lasso. In the latter case, the $\ell_{1}$-norm convergence rates are multiples of the $\ell_{2}$-norm convergence rates by a factor of the square root of the sparsity $\sqrt{s_{r}}$ or $\sqrt{s_{\pi}}$, while in the former case, the $\ell_{1}$-norm convergence rates are multiples of the $\ell_{2}$-norm convergence rates by a factor of $(\sqrt{n/\log p})^{1/(2\xi_{r}+1)}$ or $(\sqrt{n/\log p})^{1/(2\xi_{\pi}+1)}$, which can be interpreted as the square root of the “effective sparsity” in the approximately sparse models (also see Section 1 of Bradic et al. (2019a)). ∎
Finally, we state the following corollary of Theorem 3 and 5 in Section 3 that demonstrates the $\sqrt{n}$-consistency of $\hat{\tau}_{\textrm{DCal}}$ under (17). The proof sketch is deferred to Appendix B.
@@
    The statement of Theorem 3 still holds, with Condition 7 replaced by Condition 12, and $s_{r}=o(\sqrt{n}/\log p)$ replaced by $\xi_{r}>1/2$;
@@
    The statement of Theorem 5 still holds, with Condition 8 replaced by Condition 13, and $s_{\pi}=o(\sqrt{n}/\log p)$ replaced by $\xi_{\pi}>1/2$.

### v4

@@ 4.3 Extension to approximately sparse GLMs
In previous sections, we have assumed that at least one of the nuisance parameters, $\pi^{\ast}$ or $r^{\ast}$, exactly follows a sparse GLM, with sparsity below $\sqrt{n}/\log p$. Belloni et al. (2014); Chernozhukov et al. (2018); Bradic et al. (2019a) and Smucler et al. (2019) considered the so-called approximately sparse GLMs to model the nuisance parameters. Similar to Condition 3, the condition below imposes that at least one of the two nuisance parameters, $r^{\ast}$ or $\pi^{\ast}$, follows the approximately sparse GLM (note that here we adopt the definition of approximately sparse GLMs from Bradic et al. (2019a)).
Condition 10.
There exist two (nonlinear) monotonically increasing, twice-differentiable link functions $\phi,\psi$ with uniformly bounded first and second derivatives, such that either of the following holds.
  1. (i)
    The OR model can be approximated by a GLM with (nonlinear) link $\psi$ with the following properties: $p$ is sufficiently large compared to $n$ such that there exists $\beta_{p}^{\ast}$ for which
    $\sqrt{n}\left\{\mathbb{E}\left[(r^{\ast}(X)-\psi(\beta_{p}^{\ast\top}X))^{2}\right]\right\}^{1/2}=o(1),$ (18)
    and for any $s$, there exists $\xi_{r}>0$ and $\beta_{s}^{\ast}$ such that $\|\beta_{s}^{\ast}\|_{0}\leq s$ and $\|\beta_{s}^{\ast}-\beta_{p}^{\ast}\|_{2}\ll s^{-\xi_{r}}$;
  2. (ii)
    The PS model can be approximated by a GLM with (nonlinear) link $\phi$ with the following properties: $p$ is sufficiently large compared to $n$ such that there exists $\gamma_{p}^{\ast}$ for which
    $\sqrt{n}\left\{\mathbb{E}\left[(\pi^{\ast}(X)-\phi(\gamma_{p}^{\ast\top}X))^{2}\right]\right\}^{1/2}=o(1),$ (19)
    and for any $s$, there exists $\xi_{\pi}>0$ and $\gamma_{s}^{\ast}$ such that $\|\gamma_{s}^{\ast}\|_{0}\leq s$ and $\|\gamma_{s}^{\ast}-\gamma_{p}^{\ast}\|_{2}\ll s^{-\xi_{\pi}}$.
We silence the dependence on $s$ by denoting $\beta^{\ast}$ and $\gamma^{\ast}$ when $s=p$. To simplify the notation, we let $\bar{r}^{\ast}(\cdot)\coloneqq\psi(\beta^{\ast\top}\cdot)$ and $\bar{\pi}^{\ast}(\cdot)\coloneqq\phi(\gamma^{\ast\top}\cdot)$. Bradic et al. (2019a) proved the following sufficient and necessary condition for the existence of $\sqrt{n}$-consistent estimator of the ATE $\tau$ when $p$ is sufficiently large compared to $n$ (see Assumption 9, 10 and the statement of Theorem 8 of Bradic et al. (2019a), to ensure that the approximation bias in $L_{2}(\mathbb{P})$-norm by all $p$ covariates is sufficiently small):
$\xi_{r}\vee\xi_{\pi}>1/2.$ (20)
As a corollary of the results in Section 3, the DCal estimator achieves $\sqrt{n}$-consistency under (20), together with several other conditions to be stated below. We first modify the regularity conditions on the OR and PS coefficient estimates computed from the training dataset $\mathcal{D}_{\mathrm{tr}}$ as follows. Condition 6 needs to be replaced by the following:
@@
Condition 10 (i) holds true and there exists a constant $m_{\psi}$ such that for any $t$, $\psi^{\prime}(t),\psi^{\prime\prime}(t)\leq m_{\psi}$. Moreover, when $\xi_{r}>1/2$, with probability converging to $1$, there exists $\hat{\beta}$ computed from $\mathcal{D}_{\mathrm{tr}}$ such that
Similarly, Condition 7 needs to be modified as
@@
Condition 10 (ii) holds true and there exists a constant $m_{\phi}$ such that for any $t$, $\phi^{\prime}(t),\phi^{\prime\prime}(t)\leq m_{\phi}$. Moreover, when $\xi_{\pi}>1/2$, with probability converging to $1$, there exists $\hat{\gamma}$ computed from $\mathcal{D}_{\mathrm{tr}}$ such that
@@
The justification for Condition 11 and Condition 12 can be found in Appendix B.3.1. In fact, Condition 11 and Condition 12 closely resemble Condition 6 and Condition 7, respectively, which are standard results of lasso. In the latter case, the $\ell_{1}$-norm convergence rates are multiples of the $\ell_{2}$-norm convergence rates by a factor of the square root of the sparsity $\sqrt{s_{r}}$ or $\sqrt{s_{\pi}}$, while in the former case, the $\ell_{1}$-norm convergence rates are multiples of the $\ell_{2}$-norm convergence rates by a factor of $(\sqrt{n/\log p})^{1/(2\xi_{r}+1)}$ or $(\sqrt{n/\log p})^{1/(2\xi_{\pi}+1)}$, which can be interpreted as the square root of the “effective sparsity” in the approximately sparse GLMs (also see Section 1 of Bradic et al. (2019a)). ∎
Finally, we state the following corollary of Theorem 3 and 5 in Section 3 that demonstrates the $\sqrt{n}$-consistency of $\hat{\tau}_{\mathrm{DCal}}$ under (20). The proof is deferred to Appendix B.3.
@@
    The statement of Theorem 3 still holds, with Condition 6 replaced by Condition 11, and $s_{r}=o(\sqrt{n/\log p})$ replaced by $\xi_{r}>1/2$;
@@
    The statement of Theorem 5 still holds, with Condition 7 replaced by Condition 12, and $s_{\pi}=o(\sqrt{n/\log p})$ replaced by $\xi_{\pi}>1/2$.
5 Simulation studies
As discussed in Remark 6, the actual DCal approach studied theoretically in previous sections might not be computationally feasible. In this section, to alleviate the computational burden, we consider a modified DCal estimator for ATE, simply by changing the PS calibration step originally described in (10) to:
$\bm{\tilde{\pi}}\coloneqq\mathcal{L}_{\bm{\pi},\tilde{B}_{\infty}^{p}(M_{\pi})}\left(\|\cdot\|_{2}^{2};\Delta_{2,T,\pi};\mathrm{h};\mathrm{b}\right)$ (21)
with the same $\mathrm{h}$ and $\mathrm{b}$. This modified PS calibration step leads to easier implementation and faster algorithmic convergence in practice. It also mirrors closely the OR calibration step (7), in which one does not parameterize the calibrated OR estimator as GLMs. To lighten the notation, we use $\hat{\tau}_{\textrm{DCal}}$ to denote the modified DCal estimator only in this section.
Remark 10 (More comments on the modified DCal estimator).
Instead of calibrating the estimated PS regression coefficients $\hat{\gamma}$ as in the original DCal estimator, the modified DCal estimator directly calibrates the PS model. By speculating the proof of Theorem 5, the original DCal estimator, by construction of the objective function, controls $\|\hat{\gamma}-\tilde{\gamma}\|_{1}$ when $\gamma^{\ast}$ is in fact sparse, which is crucial when controlling the difference between the DCal estimator $\hat{\tau}_{\textrm{DCal}}$ and the SCal estimator $\hat{\tau}_{{\textrm{SCal},r}}$ without calibrating the PS model (when unnecessary). It is still an open question whether minimizing $\|\bm{\tilde{\pi}}\|_{2}^{2}$ directly can also imply the conclusions in Theorem 5. We emphasize that the main purpose of the simulation studies conducted here is to demonstrate that the essential idea of Double Calibration can be used to construct estimators with similar practical performance to competing methods detailed below. ∎
To study the practical merit of the modified DCal estimator, we compare it against the following competing methods: the IPW estimator $\hat{\tau}_{\mathrm{IPW}}$ (Horvitz and Thompson, 1952), the estimator $\hat{\tau}_{g}$ based on the g-formula (Robins, 1986), DML/AIPW estimator $\hat{\tau}_{\mathrm{DML}}$ (Robins et al., 1994; Hahn, 1998; Chernozhukov et al., 2018), the Targeted Maximum Likelihood Estimation (TMLE) estimator $\hat{\tau}_{\mathrm{TMLE}}$ (van der Laan and Rubin, 2006; van der Laan and Rose, 2011), the Approximate Residual Balancing (ARB) estimator $\hat{\tau}_{\mathrm{ARB}}$ (Athey et al., 2018), the Regularized Calibrated (RCal) estimator $\hat{\tau}_{\mathrm{RCAL}}$ (Tan, 2020b), and the high-dimensional Covariate Balancing Propensity Score (hdCBPS) estimator $\hat{\tau}_{\mathrm{hdCBPS}}$ (Ning et al., 2020). More details on implementing these competing methods can be found in Appendix C.1. We also set up a GitHub repository (https://github.com/Cinbo-Wang/DCal) for readers who are interested in replicating our simulation results.
We consider two different settings to mirror our theoretical results: Section 5.1 below considers sparse PS and dense OR models, whilst Section 5.2 considers the opposite setting. In both cases, the observed data is comprised of $N$ i.i.d. samples $\{X_{i},T_{i},Y_{i}\}_{i=1}^{N}$, with $X\sim\mathcal{N}_{p}(0,\Sigma)$, where $\Sigma_{i,j}\equiv 0.9^{|i-j|}$ for $i,j=1,\cdots,p$. We consider the following combinations of $(n,p)$:
$(n,p)\in\{(200,400),(400,800),(800,1000),(1600,1000)\},$
with $s\in\{10,20,50\}$ for each possible $(n,p)$ pair. For each setting, 200 Monte Carlos are sampled to get the summary statistics such as (absolute) biases, $\sqrt{N}$-biases, standard errors, mean squared errors, coverage probabilities, and etc. and all the simulated data are generated anew in each run. Variance estimators of different ATE estimators are simply based on their respective influence functions.
5.1 Case I: Sparse PS & Dense OR
First, we consider the case where the PS is a sparse logistic-linear model and the OR is a dense nonlinear model:
$T|X=x\sim\text{Bernoulli}\left(\pi(x)=\psi(x^{\top}\gamma)\right),$
$Y|X=x,T=t\sim\mathcal{N}\left(r(x,t),1\right),$
where $\psi(z)\equiv 1/(1+\exp(-z))$ in this section. In particular, we choose $\gamma$ to satisfy the following: $\|\gamma\|_{2}=1$, $\gamma_{j}\sim\mathrm{Uniform}([1,2])$ for $j=1,\cdots,s$ and $\gamma_{s+1}=\cdots=\gamma_{p}=0$ so only the first $s$ coordinates of $\gamma$ are active. For the outcome model $r(x,t)$, we first transform the covariates $X$ nonlinearly as follows:
$\tilde{X}_{1}\equiv\texttt{bs}(X_{1},100)^{\top}\cdot\left(1,1/2,1/3,\cdots,1/100\right),\tilde{X}_{2}\equiv 2/(1+\exp(-X_{2})),$
$\tilde{X}_{3}\equiv\exp(X_{3}/2),\tilde{X}_{4}\equiv X_{4}/(1+\exp(X_{3})),\tilde{X}_{5}=X_{4}X_{5}/10,$
and then scale $\tilde{X}_{2},\cdots,\tilde{X}_{5}$ to have zero mean and unit variance. Here $\texttt{bs}(x,100)$ denotes the B-spline transformation of $x$ with 100 degrees-of-freedom. We then define the outcome model $r(x,t)$ as
$r(x,t)\equiv\left(|\tilde{x}_{1}+\tilde{x}_{2}-\tilde{x}_{3}/2+\tilde{x}_{4}/3-\tilde{x}_{5}/4|+0.05\right)^{-1}+x^{\top}\beta-t,$
where $\|\beta\|_{2}\equiv 1$ and $\beta_{j}\propto j^{-1}$ for $j=1,\cdots,p$.
The simulation results are displayed in Figure 1, Table LABEL:tab:_comp_time_cont, Table LABEL:tab:_sparse_PS_measure_CI (see Appendix C.2), and Table LABEL:tab:_sparse_PS_measure (see Appendix C.2). We summarize the numerical results below:
  • Reading from root-N scaled absolute estimation errors from Figure 1, as expected, the g-formula estimator $\hat{\tau}_{g}$ generally has the largest bias as the initial OR estimator without calibration should not even be a consistent estimator. Interestingly, although the PS model is sparse, $\hat{\tau}_{\textrm{DCal}}$ still has better performance than the IPW and AIPW estimators, possibly due to its calibration step.
    $\hat{\tau}_{\mathrm{TMLE}}$, $\hat{\tau}_{\mathrm{ARB}}$, and $\hat{\tau}_{\mathrm{hdCBPS}}$ perform similarly to $\hat{\tau}_{\textrm{DCal}}$, but the performance of $\hat{\tau}_{\mathrm{TMLE}}$ deteriorates when the sparsity of the PS model is relatively large (the rightmost panel of Figure 1). $\hat{\tau}_{\mathrm{ARB}}$ generally has slightly larger bias than $\hat{\tau}_{\textrm{DCal}}$, but the difference is not very pronounced. The performance of $\hat{\tau}_{\mathrm{hdCBPS}}$ deteriorates as the sample size becomes smaller even when the sparsity of the PS model is low, but when the sample size is large, $\hat{\tau}_{\mathrm{hdCBPS}}$ performs as good as $\hat{\tau}_{\textrm{DCal}}$, and sometimes better than $\hat{\tau}_{\textrm{DCal}}$ when the sparsity of the PS model is very low (the bottom left corner of Figure 1).
    In this setting, $\hat{\tau}_{\mathrm{RCAL}}$ does not perform as well as $\hat{\tau}_{\mathrm{TMLE}}$, $\hat{\tau}_{\mathrm{ARB}}$, $\hat{\tau}_{\mathrm{hdCBPS}}$ and $\hat{\tau}_{\textrm{DCal}}$; we conjecture that this is due to the relatively low overlap in propensity score and in Appendix C.2 we indeed show that $\hat{\tau}_{\mathrm{RCAL}}$ has much more improved performance when the violation of the positivity assumption is alleviated. Similar conclusion can be reached by reading from Tables LABEL:tab:_sparse_PS_measure_CI and LABEL:tab:_sparse_PS_measure, which respectively report more detailed numerical values of the Mean/Median Absolute Biases (MAB), standard errors, and Root-Mean-Squared-Errors (RMSE), and the coverage probabilities and lengths of 95% Wald confidence intervals centered around different ATE estimators.
  • Reading from the left panel of Table LABEL:tab:_comp_time_cont, it is quite obvious that $\hat{\tau}_{\textrm{DCal}}$ enjoys shorter computation time than $\hat{\tau}_{\mathrm{hdCBPS}}$ and $\hat{\tau}_{\mathrm{RCAL}}$.
Figure 1: Boxplots of the $\sqrt{n}$-scaled estimation error $\sqrt{n}|\hat{\tau}-\tau|$ under different sample size $n$, dimension $p$ and sparsity level $s$ under the Sparse PS & Dense OR setting. The white dots correspond to mean.
5.2 Case II: Dense PS & Sparse OR
Next, we consider the case where the PS is a dense nonlinear model bounded between 0 and 1 and the OR is a sparse linear model:
$T|X=x\sim\mathrm{Bernoulli}\left(\pi(x)\right),$
$Y|X=x,T=t\sim\mathcal{N}\left(r(x,t),1\right).$
Here we take the PS model to be
$\pi(X)\equiv\min\{\max\{\psi(\check{X}_{1}-\frac{1}{2}\check{X}_{2}+\frac{1}{4}\check{X}_{3}-\frac{1}{8}\check{X}_{4}+\gamma^{\top}X),0.05\},0.95\},$
where $\check{X}_{1}\coloneqq e^{0.5X_{1}}$, $\check{X}_{2}\coloneqq 10+X_{2}/(1+e^{X_{1}})$, $\check{X}_{3}\coloneqq(0.05X_{1}X_{3}+0.6)^{2}$, $\check{X}_{4}\coloneqq(X_{2}+X_{4}+10)^{2}$, and $\gamma_{j}\propto j^{-1}$ with $\|\gamma\|_{2}\equiv 2$. We take the OR model to be
$r(X,t)\equiv\beta^{\top}X-1+2t$
where we first draw $\beta_{j}\sim\mathrm{Uniform}([1,2])$ for $j=1,\cdots,s$ and set $\beta_{j}\equiv 0$ for $j=s+1,\cdots,p$. We eventually normalize $\beta$ such that $\|\beta\|_{2}=1$.
The simulation results are displayed in Figure 2, Table LABEL:tab:_comp_time_cont, Table LABEL:tab:_sparse_OR_measure_CI (see Appendix C.3), and Table LABEL:tab:_sparse_OR_measure (see Appendix C.3). We summarize the numerical results below:
  • Reading from root-N scaled absolute estimation errors from Figure 2, as expected, the IPW estimator $\hat{\tau}_{\mathrm{IPW}}$ generally has the worst bias as the initial PS estimator without calibration is not even a consistent estimator. Interestingly, although the OR model is sparse, $\hat{\tau}_{\textrm{DCal}}$ still has better performance than the g-formula and AIPW estimators, possibly due to its calibration step.
    $\hat{\tau}_{\mathrm{TMLE}}$, $\hat{\tau}_{\mathrm{ARB}}$, and $\hat{\tau}_{\mathrm{hdCBPS}}$ perform similarly to $\hat{\tau}_{\textrm{DCal}}$, but the performance of $\hat{\tau}_{\mathrm{TMLE}}$ again deteriorates when the sparsity of the OR model is relatively large (the rightmost panel of Figure 2). In this case, $\hat{\tau}_{\mathrm{ARB}}$ and $\hat{\tau}_{\mathrm{hdCBPS}}$ perform as well as $\hat{\tau}_{\textrm{DCal}}$. This is not very surprising at least for $\hat{\tau}_{\mathrm{ARB}}$, since it is proved that $\hat{\tau}_{\mathrm{ARB}}$ is $\sqrt{n}$-consistent in this regime.
    $\hat{\tau}_{\mathrm{RCAL}}$ does not perform as well as $\hat{\tau}_{\mathrm{ARB}}$, $\hat{\tau}_{\mathrm{hdCBPS}}$ and $\hat{\tau}_{\textrm{DCal}}$: the performance of $\hat{\tau}_{\mathrm{TMLE}}$ recovers when the sample size increases, whereas for the performance of $\hat{\tau}_{\mathrm{RCAL}}$ we again conjecture that this is due to the relatively large $\ell_{2}$-norm of $\gamma$ (recall that $\|\gamma\|_{2}=2$) and in Appendix C.3 we indeed show that $\hat{\tau}_{\mathrm{RCAL}}$ has much more improved performance when $\|\gamma\|_{2}$ decreases. It is worth noting that by looking at the results across different sample sizes, $\hat{\tau}_{\mathrm{RCAL}}$ is still close to being $\sqrt{n}$-consistent empirically and the main difference between $\hat{\tau}_{\mathrm{RCAL}}$ and other estimators is in the constant.
    Similar conclusion can be reached by reading from Tables LABEL:tab:_sparse_OR_measure_CI and LABEL:tab:_sparse_OR_measure, which respectively report the coverage probabilities and lengths of 95% Wald confidence intervals centered around different ATE estimators, and more detailed numerical values of the Mean/Median Absolute Biases (MAB), standard errors, and Root-Mean-Squared-Errors (RMSE).
  • Reading from the right panel of Table LABEL:tab:_comp_time_cont, it is quite obvious that $\hat{\tau}_{\textrm{DCal}}$ again enjoys much shorter computation time than $\hat{\tau}_{\mathrm{hdCBPS}}$ and $\hat{\tau}_{\mathrm{RCAL}}$.
Figure 2: Boxplots of the $\sqrt{n}$-scaled estimation error $\sqrt{n}|\hat{\tau}-\tau|$ under different sample size $n$, dimension $p$ and sparsity level $s$ under the Dense PS & Sparse OR setting. The white dots correspond to means.
Table 2: Computation time (in minutes) of different methods averaged over 200 Monte Carlos.
$n$ $p$ $s$ Sparse PS & Dense OR Dense PS & Sparse OR
[…]
200 400 10 0.08 0.62 0.16 0.02 1.40 0.45 0.11 0.02
200 400 20 0.10 0.59 0.15 0.02 2.06 0.46 0.12 0.02
200 400 50 0.06 0.45 0.11 0.02 2.08 0.47 0.12 0.02
400 800 10 0.10 4.33 0.97 0.06 3.10 2.71 0.55 0.07
400 800 20 0.20 5.15 1.01 0.06 4.52 2.67 0.51 0.07
400 800 50 0.05 4.81 0.88 0.06 4.55 2.67 0.51 0.08
800 1000 10 0.64 60.77 6.84 0.16 7.47 24.41 2.72 0.19
800 1000 20 0.30 43.52 7.18 0.16 10.19 24.30 2.36 0.21
800 1000 50 0.19 48.06 6.01 0.16 10.09 24.24 2.26 0.18
1600 1000 10 3.37 437.60 22.48 0.36 22.58 254.20 5.99 0.44
1600 1000 20 1.87 513.17 24.26 0.34 27.20 248.58 5.03 0.41
1600 1000 50 0.62 328.59 26.79 0.34 27.19 244.39 4.58 0.43